OpenStudy (anonymous):

8th grade math help please

OpenStudy (anonymous):

roy can order a large or small snow cone. There are 15 flavors and 2 toppings from which to choose. How many ways can Troy order a snow cone? A. 32 B. 60 C. 120 D. 30

OpenStudy (anonymous):

It's D)30

OpenStudy (anonymous):

multiply 15 by 2. @i_love_bands26 , do not give the answer. you are supposed to help them solve it by themselves.

OpenStudy (anonymous):

A cooking club has 12 members. How many ways can they select a president and vice president from among the members? A. 1320 B. 66 C. 132 D. 220

OpenStudy (anonymous):

You guys forgot the toppings. Let 'S' stand for sizes, 'F' for flavors, 'T' for toppings, and 'C' for total combinations. It should go as follows: (S x F)(S x T)(F x T)= C (2 x 15)(2 x 2)(15 x 2) = C At least, I think that's it.

OpenStudy (anonymous):

Er, (2 x 15) + (2 x 2) + (15 x 2) = C.

OpenStudy (anonymous):

so is 30 right? I need to pass this

OpenStudy (anonymous):

No. It's not thirty.

OpenStudy (anonymous):

Try this out, it should be it: (2 x 15) + (2 x 2) + (15 x 2) = C.

OpenStudy (anonymous):

oh the first question is B)60